The Sisyphean Corporation is considering investing in a new cane manufacturing machine with a cost of $30,000. It is to be depreciated in a continuing pool with a CCA rate of 50 percent. The machine has no resale value and will produce sales of 2,000 in year 1. Sales are estimated to grow by 10% per year each year through year three. The price per cane that Sisyphean will charge its customers is $18 each and is to remain constant. The canes have a cost per unit to manufacture of $9 each. Installation of the machine and the resulting increase in manufacturing capacity will require an increase in various net working capital accounts. It is estimated that the Sisyphean Corporation needs to hold 2% of its annual sales in cash, 4% of its annual sales in accounts receivable, 9% of its annual sales in inventory, and 5% of its annual sales in accounts payable. The firm is in the 35% tax bracket and has a cost of capital of 10%. The incremental EBIT in the first year for the Sisyphean Corporations project is closest to: $7,500 $6,825 $10.500 $18,000
Answer #1

As per given information,

In Year 1, Sales = 2000 units

Selling Price per units = 18

Sales Revenue = 2000 * 18 = 36000

Manufacturing Cost Per unit = 9

Manufacturing Cost = 2000*9 = 18000

EBIT = Sales Revenue - Manufacturing Cost = 36000 - 18000

EBIT = 18000